OpenStudy (saifoo.khan):

Completing square (Show Steps): f(x) = -0.6 x^2+1.2 x+1.4 What is the Domain & Range?

OpenStudy (anonymous):

what is with you with these problems???

OpenStudy (anonymous):

lets to it the easy way \[-\frac{b}{2a}=-\frac{1.2}{2\times -.6}=2\] so you will have \[f(x)=-.6(x-2)^2+\text{something}\] and that "something" is \[f(2)\]

OpenStudy (saifoo.khan):

f(x) = -(0.6) (x - 1)² + 2

OpenStudy (saifoo.khan):

that's the answer/

OpenStudy (anonymous):

oh cramp that is wrong! it is \[-\frac{b}{2a}=-\frac{1.2}{2\times -.6}=1\]

OpenStudy (saifoo.khan):

I didn't liked yesterday one, so made a new one! ;)

OpenStudy (anonymous):

yeah 1 ! my arithmetic is bad

OpenStudy (anonymous):

what is this ? and saifoo you asked a question?! i have only seen you answer them

OpenStudy (anonymous):

so it will be \[f(x)=-0.6(x-1)^2+\text { whatever } f(1) \text{ is}\]

OpenStudy (saifoo.khan):

haha violinplayer, i suck @ completing square.. so i asked for help! ;)

OpenStudy (anonymous):

oh

OpenStudy (saifoo.khan):

-0.6 x^2+1.2 x+1.4 \[-0.6(-1)^2 +1.2(-1) + 1.4\] Like this?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Did you just ask and answer your own problem?

OpenStudy (saifoo.khan):

haha, yea.